9. HF has a Ka = 6.8*10. Find the pH of a 0.15M solution of its % - ionization. HF and
Answer #1

Aso ko [F] [H0] [HE] xox Rx 10 109 = (0.15-2) 6.8 x 10 x = x - 1.02xlo - - 1.02x10 20 -) x + 6.8 x 104 x by soling this quaCrinen! [HE] = 0.15 M ka= 6.8 8109 — Е- + но* НЕ + Но 0.15M то у ) О, 15 – х. Ics Scanned with CamScanner

Hope it helps...pls do upvote